# AI-Driven Skin Cancer Diagnosis: Grad-CAM and Expert Annotations for Enhanced Interpretability

### v1

@@ Abstract
An AI tool has been developed to provide interpretable support for the diagnosis of BCC via teledermatology, thus speeding up referrals and optimizing resource utilization. The interpretability is provided in two ways: on the one hand, the main BCC dermoscopic patterns are found in the image to justify the BCC/Non BCC classification. Secondly, based on the common visual XAI Grad-CAM, a clinically inspired visual explanation is developed where the relevant features for diagnosis are located. Since there is no established ground truth for BCC dermoscopic features, a standard reference is inferred from the diagnosis of four dermatologists using an Expectation Maximization (EM) based algorithm. The results demonstrate significant improvements in classification accuracy and interpretability, positioning this approach as a valuable tool for early BCC detection and referral to dermatologists. The BCC/non-BCC classification achieved an accuracy rate of 90%. For clinically-inspired XAI results, the detection of BCC patterns useful to clinicians reaches 99% accuracy. As for the Clinically-inspired Visual XAI results, the mean of the Grad-CAM normalized value within the manually segmented clinical features is 0.57, while outside this region it is 0.16. This indicates that the model struggles to accurately identify the regions of the BCC patterns. These results prove the ability of the AI tool to provide a useful explanation.
Keywords Artificial Intelligence  $\cdot$ Basal Cell Carcinoma  $\cdot$ Deep Learning  $\cdot$ Skin Cancer  $\cdot$ Telemedicine  $\cdot$ XAI
@@ 1 Introduction
Skin cancer is the most commonly diagnosed cancer worldwide, with Melanoma, Basal Cell Carcinoma (BCC), and Squamous Cell Carcinoma (SCC) being the most prevalent types. BCC accounts for approximately $75$% of all skin cancers, making it the most frequent form. It has well-established clinical criteria for diagnosis, yet there is significant variability in the presence of these clinical criteria between cases [1, 2, 3, 4].
@@
Many research efforts seek to diagnose various skin diseases. Additionally, the number of published papers has increased significantly in recent years, driven by the development of public databases [5, 6, 7, 8]. Although these databases are accessible and comprehensive for everyone, the clinical criteria used for diagnosis are not readily available. To develop a tool that is useful from a medical perspective, it is crucial to provide not only classification metrics, but also a detailed diagnosis explaining the detected clinical features, thereby offering a more comprehensive result. In this context, Serrano at al. [9] developed a clinically inspired skin lesion classification tool through the detection of dermoscopic criteria for BCC.
@@
XAI techniques such as Grad-CAM [10] play a crucial role in improving the transparency and interpretability of CNNs in any kind of medical diagnostics. By offering visual explanations of the decision-making process, these methodologies allow medical professionals to gain insight into the predictive mechanisms of AI models, thus fostering trust and acceptance in AI-driven diagnostics. In recent years, XAI has become an active field of research and numerous review papers have been published [11, 12, 13].
In particular, there are some recent papers on XAI in skin image analysis. Barata et al. uses attention modules to improve explainability in hierarchical skin lesion diagnosis [14]. Chanda et al. developed a multimodal XAI system to support dermatologists in diagnosing melanoma, which offers text and region-based explanations alongside its diagnostic predictions [15]. Rezk et al. developed an interpretable skin cancer diagnosis model using clinical images to assist general practitioners in early detection and referral. Their approach incorporates visual explanations of the model’s decisions based on Grad-CAM++ [16].
In these previous works, XAI is used to visualize the regions where the models have focused to extract features. The proposed method goes a step further: It proposes a clinical interpretation by localizing regions of clinical interest for diagnosis, specifically BCC dermoscopic features, and providing clinical labels.
@@ 2 Methodology
In modern healthcare, primary care physicians use teledermatology systems to receive high-quality diagnostic images remotely, allowing preliminary diagnoses of Basal Cell Carcinoma (BCC) using established patterns [4, 17]. Suspected BCC cases are promptly referred to dermatology specialists, improving healthcare efficiency, reducing waiting times, and facilitating early intervention.
Teledermatology is crucial in areas with limited specialist access and its integration with AI tools promises to enhance diagnostic accuracy and efficiency. This work aims to develop an AI tool to assist in this process by providing a binary classification of BCC/non-BCC with interpretable results Fig. 1 illustrates the current and proposed diagnostic workflows.
Figure 1: Illustration showing the current workflow of BCC diagnosis by teledermatology in orange and the proposed workflow in green.
@@ 2.1 Database
The entire database was provided by the Dermatology Unit of the “Hospital Universitario Virgen Macarena” and were sent over 2 years from 60 primary care centers. The dataset comprises 1559 dermoscopic images divided into 3 subsets. Four dermatologists provided different types of annotation according to the subsets. Specifically:
  • The first subset consisted of 1089 dermoscopic images. Initially, the labeling annotations for these images were the presence or absence of each of the dermoscopic features involved in the diagnosis of BCC.
  • A second subset of 334 images is additionally enriched with dermatologist delineations of BCC dermoscopic patterns within each image. More than one segmented area may appear on an image if there are multiple patterns in the BCC lesion. In the Fig. 3 an example is shown.
  • The third subset is made up of 136 non-BCC images, mostly consisting of nevus lesions, from the ISIC archive [8].
Table 1 summarizes the distribution of labels in the database. As can be seen in this table, the database has a significant class imbalance, with SW and MG underrepresented. Several techniques have been used to address this problem.
@@
2.1.1 Label codification
Each image may contain multiple dermoscopic patterns. Therefore, a one-hot coding scheme was used to encode the labels during image annotation and subsequently to process the dermatologists’ annotations. Each image label is a binary word and each BCC dermoscopic pattern is a digit, where $1$ means presence and 00 means absence. The seven patterns that can appear in a BCC lesion are[4, 18, 17]: Pigment Network (PN) (negative criterion), Ulceration (U), Ovoid Nests (ON), Multiglobules (MG), Maple Leaf-like (ML), Spoke Wheel (SW), Arborizing Telangiectasia (AT) (Fig. 2). Thus, each label is a vector of dimensions $[1\textrm{x}7]$. In Table 2 there are some examples of this process.
2.1.2 Standard Reference Inferring
The accepted ground truth (GT) for BCC diagnosis is biopsy. However, there is no established GT for BCC dermoscopic patterns, which are subjectively assessed by dermatologists. Several studies have reported a low kappa coefficient when measuring inter-dermatologist agreement in determining the different dermoscopic patterns present in a lesion [19, 17]. Therefore, an adequate SR inferred from the consensus of several dermatologists is required. An Expectation-Maximization (EM) based algorithm [20] was implemented to derive a single SR for model training from multiple specialist labels. This algorithm consolidates multilabel annotations from different dermatologists and generates a single inferred SR that encapsulates the collective expertise of the raters. Silva et al. [21] used this algorithm to infer and SR from BCC pattern annotations and demonstrated that integrating this diverse expertise mitigates the subjectivity inherent in diagnosing the BCC pattern and improves the reliability and robustness of the classification model.
Table 2: Example of multilabel and binary encoding for BCC diagnosis
Codification Multi-label Binary Diagnostic
Example 1 [0 1 0 1 1 0 1] 1 Presence of BCC
Example 2 [1 0 0 0 0 0 0] 0 Absence of BCC
Example 3 [0 0 0 0 0 0 0] 0 Absence of BCC
2.2 Clinical inspired XAI
@@
From a clinical point of view, achieving 100% accuracy in detecting BCC dermoscopic patterns is not crucial for a correct diagnosis of BCC. A dermatologist requires detecting just one correct BCC dermoscopic pattern to make an accurate BCC diagnosis. Therefore, a useful clinically-inspired XAI should provide any of the following outcomes: if no patterns are detected, it predicts non-BCC; if the PN pattern is detected, it serves as a negative criterion, explaining a non-BCC prediction; and if any other BCC pattern is detected, the prediction is positive for BCC.
@@
The proposed AI tool for skin lesion diagnosis is based on the MobileNet-V2 model with 3 classifier layers. This model was joined with an optimization and training strategy in three phases. In the first step, ImageNet transfer learning was applied, and the classifier weights were trained. In the second step, fine-tuning was applied to the last three blocks of the feature extractor and the classifier with a lower learning rate (LR) and increasing the number of epochs. In the third step, transfer learning was applied from the binary model to solve the task of BCC pattern detection, retraining only the classifier with a very low LR
2.3 Clinical visual XAI
@@
To develop this concept, expert-generated segmentations are used. As detailed in Sect. 2.1, a subset of samples was segmented by a specialist. A dermatologist performed an individual segmentation of each BCC pattern for each image.
@@
Understanding the decision-making process of convolutional neural networks (CNNs) is critical for clinical applications where interpretability is as important as performance metrics. The most common visualization techniques in explainable AI (XAI) are activation maps and gradient-weighted class activation mapping (Grad-CAM). Activation Maps, as described by Zhou et al. [22], extract features from different network layers to show which patterns the model focuses on, although they may not always directly correlate with specific outcomes. On the other hand, Grad-CAM, as described in detail by Selvaraju et al. [10], uses the gradients of the predicted class that flow into the final convolutional layer to produce a localization map. The clinically-inspired explanation that we proposed in this paper is based on Grad-CAM.
@@
However, additional information is provided. For this study, the segmentations of individual BCC patterns (see 2.1) were combined into a single segmented image, as shown in Fig. 3. And these manual segmentations were overlaid with activation maps to validate the clinical information provided by Grad-CAM. In this way, the explanation is not only where the AI tool is paying attention, but also what the dermoscopic pattern of that region is.
(a) BCC lesion
(b) All patterns
(c) Maple Leaf
(d) Ovoid Nests
(e) Multiglobules
Figure 3: Example of BCC dermoscopic patterns segmentation by a dermatologist
@@ 3.1 Implementation details
The chosen optimizer was the AdamW schedule-free optimizer [23], with a mini-batch size of 32, and dropout rate with a rate of $0.3$ to prevent overfitting. Focal loss [24] has been used to address class imbalance.
Data augmentation techniques [25, 26] applied included rotation, perspective transformation, and Gaussian blur.
Due to the limited size of our database, a stratified k-fold cross-validation was implemented to ensure a comprehensive evaluation of the model’s performance. This approach mitigates the risk of biased training and testing distributions, especially crucial in datasets with imbalanced class distributions.
3.2 Clinical-inspired XAI: BCC diagnosis with additional label information

### v2

# Clinically inspired enhance Explainability and Interpretability of an AI-Tool for BCC diagnosis based on expert annotation

@@ Abstract
Background and Objective: An AI tool has been developed to provide interpretable support for the diagnosis of BCC by teledermatology, thus speeding up referrals and optimizing resource utilization. Dermatologists base their BCC diagnosis on the presence or absence of specific BCC dermoscopic features. Our clinically inspired explainability approach relies on two key pieces of evidence: the BCC dermoscopic patterns identified in the image by the AI classification system and the image regions where the AI tool focuses, which correspond to the presence of BCC patterns.
Methods: A dataset of 1559 dermoscopic images was annotated by several dermatologists. MobileNet-V2 was employed for BCC/non-BCC classification and BCC pattern detection, developing a clinically-inspired XAI approach. The model was trained using stratified cross-validation and data augmentation to address class imbalance. Visual XAI was implemented to validate the clinically-inspired XAI detection. Since there is no established ground truth for BCC dermoscopic features, a standard reference was inferred from the diagnoses of four dermatologists using an algorithm based on expectation maximization (EM).
Results: The results demonstrate significant improvements in classification accuracy and interpretability, positioning this approach as a valuable tool for early BCC detection and referral to dermatologists. The BCC/non-BCC classification achieved an accuracy rate of 90%. For clinically-inspired explainable AI (XAI) results, the detection of BCC patterns useful to clinicians reaches 99% accuracy. As for the Clinically-inspired Visual XAI results, the mean of the Grad-CAM normalized value within the manually segmented clinical features is 0.57, while outside this region it is 0.16. This indicates that the model struggles to accurately identify the regions of the BCC patterns. These results demonstrate the ability of the AI tool to provide a useful explanation.
Conclusions: Based on these results, two significant clinical achievements have been attended: (1) A tool with dual explanation capabilities rooted in clinical practice has been developed, which improves reliability and adoption by dermatologists. (2) This advancement allows integration into the clinical field, which could change the current protocol in primary care centers to accelerate treatment for patients with BCC lesions in various areas of Andalusia.
keywords:
Basal Cell Carcinoma , Deep Learning , Skin Cancer , Telemedicine , XAI
journal: Artificial Intelligence in Medicine
[…]
[1]organization=University of Seville, city=Seville, country=Spain
[…]
[2]organization=Hospital Universitario Virgen de Macarena, city=Seville, country=Spain
[…]
[3]organization=Hospital Quirón Salud Infanta Luisa y Sagrado Corazón, city=Seville, country=Spain
@@ 1 Introduction
Worldwide, doctors diagnose skin cancer most commonly, with Melanoma, Basal Cell Carcinoma (BCC), and Squamous Cell Carcinoma (SCC) as the most prevalent types. BCC accounts for approximately $75$% of all skin cancers, making it the most frequent form. It has well-established clinical criteria for its diagnosis, yet there is significant variability in the presence of these clinical criteria between cases [1, 2, 3, 4].
@@
The development of public databases has significantly increased the number of published articles related to automatic detection of skin diseases in recent years [5, 6, 7, 8]. Although these databases are accessible and comprehensive for everyone, the clinical criteria used for diagnosis are not readily available. To develop a tool that is useful from a medical perspective, it is crucial to provide not only classification metrics, but also a detailed diagnosis explaining the detected clinical features, offering a more comprehensive result. In this context, Serrano et al. [9] created a clinically inspired skin lesion classification tool by detecting dermoscopic criteria for BCC.
@@
XAI techniques such as Grad-CAM [10] play a crucial role in improving the transparency and interpretability of CNNs in any kind of medical diagnosis. By offering visual explanations of the decision-making process, these methodologies allow medical professionals to gain insight into the predictive mechanisms of AI models, thus fostering trust and acceptance in AI-driven diagnoses. The integration of XAI with dermatological applications not only enhances the model’s interpretability but also aids in the early and accurate diagnosis of skin conditions, potentially leading to better patient outcomes. In van der Valden et al. survey [11] there are several works focused on the diagnosis of skin cancer that use different XAI techniques to see how their models behave are explained.
Barata et al. have developed a Computer-Aided Diagnosis (CAD) system based on deep learning for the hierarchical diagnosis of skin lesions, using ISIC 2017 and 2018 datasets. Their model uses attention modules to improve explainability by highlighting important features in dermoscopy images. This approach improves the transparency and interpretability of the diagnostic process, making the system more reliable and suitable for clinical use [12].
Chanda et al. developed a multimodal XAI system to support dermatologists in diagnosing melanoma. Using the HAM10000 dataset, which comprises over 10,000 dermoscopic images, they created a model that provides interpretable, text and region-based explanations alongside its diagnostic predictions. A three-phase reader study involving 116 international dermatologists demonstrated that the XAI system significantly increased clinicians’ trust and confidence in their diagnosis compared to a conventional AI system, although it did not significantly improve diagnostic accuracy [13].
Rezk et al. developed an interpretable skin cancer diagnosis model using clinical images to assist general practitioners in early detection and referral. Their approach incorporates a skin lesion taxonomy into a deep neural network to incrementally learn dermatological concepts. They used a dataset of 914 clinical images, supplemented by additional images from DermNet NZ to address class imbalances. To enhance transparency and trustworthiness in line with recent regulatory requirements, XAI techniques, such as Grad-CAM++, were employed to provide visual explanations of the model’s decisions. The proposed model significantly improved classification accuracy and interpretability, facilitating better understanding and detection of misdiagnoses [14].
However, in [12, 14] works, the use of XAI is employed to visualize the regions where the models have focused to extract features, providing a visual representation of the results returned by the classifier. Besides, [13] introduces a novel use of Grad-CAM to improve dermatologist trust by ensuring the AI model focuses on clinically relevant lesion areas. Grad-CAM explanations are combined with feature detection, aligning AI decisions with expert reasoning, significantly increasing user confidence without sacrificing diagnostic accuracy. Nevertheless, the key contribution of this work lies in a dual-layer explanation aimed at enhancing clinical diagnosis. Firstly, it offers a binary detection between BCC and non-BCC cases. Secondly, it provides a clinically valuable XAI tool for specialists, enabling them to verify whether its attention is being directed towards the correct BCC-related patterns.
@@ 2 Methodology
In modern healthcare, primary care physicians use teledermatology to receive high-quality diagnostic images remotely, allowing preliminary diagnoses of Basal Cell Carcinoma (BCC) using established patterns [4, 15]. Suspected BCC cases are immediately referred to dermatology specialists, improving healthcare efficiency, reducing waiting times, and facilitating early intervention.
Teledermatology is crucial in areas with limited specialist access and its integration with AI tools promises to enhance diagnostic accuracy and efficiency. This work aims to develop an AI tool to assist in this process by providing a binary classification of BCC/non-BCC with interpretable results Fig. 1 illustrates the current and proposed diagnostic workflows. In the proposed workflow, the AI tool not only delivers a BCC/non-BCC diagnosis but also provides a layer of interpretability, guiding the dermatologist to focus on the dermoscopic features most relevant to basal cell carcinoma. This dual-layer explanation ensures that the specialist can validate the AI’s predictions by confirming that the model is emphasising the correct clinical patterns, thus reducing the chances of diagnostic errors and increasing trust in AI-assisted teledermatology.
Figure 1: Illustration showing an example of the current workflow of BCC diagnosis by teledermatology in orange and the proposed workflow in green.
@@ 2.1 Database
The Dermatology Unit of the “Hospital Universitario Virgen Macarena” provided the entire database, which consists of dermoscopic images sent from 60 primary care centers during 2 years. The dataset comprises 1559 dermoscopic images divided into 3 subsets. Four dermatologists provided different types of annotation according to the subsets. Specifically:
  • 1.
    The first subset consists of 1089 dermoscopic images that were annotated following a multilabel codification explained in Sect. 2.1.1. To consolidate the multi-rater labeling, an Expectation-Maximization (EM) algorithm was employed, generating an inferred Standard Reference (SR), which will be explained in Sect. 2.1.2. Subsequently, a binary classification scheme was established to classify into BCC and non-BCC lesions.
  • 2.
    An additional set of 334 images was collected, enriched with detailed dermatological annotations beyond the multilabel and binary SR. It includes specialist’s manual segmentations of the BCC clinical patterns present in a lesion. Several segmented area may appear on an image if there are some patterns in the BCC lesion. In the Fig. 3 an example is shown. These images are designated for the XAI part of the study, enabling a deeper analysis of the model decision process.
  • 3.
    The final subset was gathered from the ISIC Archive [8], utilizing 136 non-BCC images, predominantly consisting of nevus lesions.
2.1.1 Label codification
Each image may contain multiple dermoscopic patterns. Therefore, we used a one-hot coding scheme to encode the labels during image annotation and subsequently process the dermatologists’ annotations. Each image label is a binary word and each BCC dermoscopic pattern is a digit, where $1$ means presence and 00 means absence. The seven patterns (Fig. 2) that can appear in a BCC lesion are[4]: Pigment Network (PN) (negative criterion), Ulceration (U), Ovoid Nests (ON), Multiglobules (MG), Maple Leaf-like (ML), Spoke Wheel (SW), Arborizing Telangiectasia (AT). Thus, each label is a vector of dimensions $[1\textrm{x}7]$ using a binary codification where a value of 1 means BCC lesion and a value of 0 means non-BCC lesion. In Table 1 there are some examples of this process.
2.1.2 Standard Reference (SR) Inferring
The accepted ground truth (GT) for BCC diagnosis is biopsy. However, there is no established GT for BCC dermoscopic patterns, which are subjectively assessed by dermatologists. Several studies have reported a low kappa coefficient when measuring inter-dermatologist agreement in determining the different dermoscopic patterns present in a lesion [16, 15]. Therefore, an adequate SR inferred from the consensus of several dermatologists is required. We implemented an Expectation-Maximization (EM) based algorithm [17] to derive a single SR for model training from multiple specialist labels. This algorithm consolidates multilabel annotations from different dermatologists and generates a single inferred SR that encapsulates the collective expertise of the raters. Silva et al. [18] used this algorithm to infer the SR from BCC pattern annotations and demonstrated that integrating this diverse expertise mitigates the subjectivity inherent in diagnosing the BCC pattern and improves the reliability and robustness of the classification model.
Table 1: Example of multilabel and binary encoding for BCC diagnosis
Codification Multi-label Binary Diagnosis
Example 1 [0 1 0 1 1 0 1] 1 Presence of BCC
Example 2 [0 0 1 0 0 0 0] 1 Presence of BCC
Example 2 [1 0 0 0 0 0 0] 0 Absence of BCC
Example 3 [0 0 0 0 0 0 0] 0 Absence of BCC
@@ 2.1.2 Standard Reference (SR) Inferring
Table 2 summarises the distribution of labels in the database. As can be seen in this table, the database has a significant class imbalance, with SW and MG patterns under-represented. Several techniques have been used to address this problem, see Sect. 3.1.
2.2 Clinically-inspired XAI: BCC diagnosis with additional label information
@@
This section refers to the AI model positioned at the bottom of Novel telemedicine diagnosis in Fig. 1 (Pattern detection AI). From a clinical point of view, achieving 100% accuracy in the detection of dermoscopic BCC patterns is not crucial to a correct diagnosis of BCC. A dermatologist only needs to detect one of the aforementioned BCC patterns to decide whether the lesion is BCC. Therefore, a useful XAI should present the same criterion. The proposed tool presents the following results: if no pattern is detected, it predicts no BCC; if the PN pattern is detected, it serves as a negative criterion, obtaining a prediction of no BCC; and if any other BCC pattern is detected, the prediction is positive for BCC.
@@
The proposed AI tool is based on the MobileNet-V2 model with three classifier layers. The model was coupled with a three-step optimization and training strategy. The first two steps focused on the detection of BCC/non-BCC, while the third step focused on pattern classification. In the first step, ImageNet transfer learning was applied and classifier weights were trained. In the second step, fine-tuning was applied to the last three blocks of the feature extractor, and the classifier was tuned with a lower learning rate (LR) and an increased number of epochs. In the third step, once the column extracts features from the same region as the specialist in the binary task, the classifier was retrained for pattern detection with a very low LR.
In light of the considerable class imbalance evident in Table 2, the methodology was devised to circumvent a direct transition from the adjusted ImageNet weights to patterns characterized by such an imbalance. Instead, an intermediate phase was incorporated, during which the weights were fine-tuned on the dermoscopic images. This intermediate phase allowed the extraction of features specific to dermoscopic images, thereby enhancing the subsequent classification process and ensuring greater accuracy and manageability.
2.3 Clinically-inspired Visual XAI
@@
This section refers to the AI model at the top of Novel telemedicine diagnosis in Fig. 1 (BCC/Non-BCC classification). To develop this concept, expert-generated segmentations are used. As detailed in Sect. 2.1, a subset of samples was segmented by a specialist. A dermatologist performed an individual segmentation of each BCC pattern for each image (Fig. 3).
@@
Understanding the decision-making process of convolutional neural networks (CNNs) is critical for clinical applications where interpretability is as important as performance metrics. The most common visualization techniques in XAI are activation maps and gradient-weighted class activation mapping (Grad-CAM). Activation Maps, as described by Zhou et al. [19], extract features from different network layers to show which patterns the model focuses on, although they may not always directly correlate with specific outcomes. On the other hand, Grad-CAM, as described in detail by Selvaraju et al. [10], uses the gradients of the predicted class that flow into the final convolutional layer to produce a localization map. The clinically-inspired explanation that we proposed in this paper is based on Grad-CAM.
@@
However, additional information for validation is provided. For this study, the segmentations of individual BCC patterns (see 2.1) were combined into a single segmented image, as shown in Fig. 3. The manual segmentations were then overlayed with activation maps to validate the clinical information provided by Grad-CAM. This not only confirmed whether the network is correctly trained in BCC/non-BCC detection, allowing us to use that spine for pattern detection as previously explained, but also identified the dermoscopic pattern of that region verifying if indeed clinical XAI is working correctly.
(a) Image example of specialist’s segmentation.
(b) Full lesion segmentation with all pattern segmentations.
(c) Maple Leaf-like pattern present on the original image.
(d) Second Maple Leaf-like pattern present on the original image.
(e) Ovoid Nests pattern present on the original image.
(f) Multiglobules pattern present on the original image.
Figure 3: Example of the specialist segmentation job.
@@ 3.1 Implementation details
Optimisation of the model as a key element of this project. Several critical hyperparameters were carefully adjusted in order to improve performance. The network architecture was trained in two stages. First, ImageNet’s transfer learning stage used the new AdamW schedule-free optimiser [20] with a learning rate of $1e^{-5}$ for 100 epochs, employing a mini-batch size of 32. A dropout rate of $0.3$ was established to serve as a regularisation technique to prevent overfitting.
As demonstrated in Sect. 2.1, class imbalance is a significant issue, especially in the multi-label classification of BCC patterns such as ML, MG, or SW. To address this problem, the Focal loss [21] was implemented, which is particularly useful when the distribution of samples is not uniform.
Due to the restricted scope of our database, we have employed a stratified k-fold cross-validation to guarantee a thorough assessment of the model’s efficacy. Specifically, a $5$-fold stratification [22, 23] was used, ensuring that each fold was well-representative of the whole dataset, maintaining the proportion of each class in every fold. This approach mitigates the risk of biased training and testing distributions, which is crucial in datasets with imbalanced class distributions.
For the classification model, we chose MobileNet-V2222, a lightweight deep neural network known for its efficiency and effectiveness in image classification tasks. To enhance the robustness of the model and account for variabilities in real-world dermoscopic imaging, several data augmentation techniques were applied [24, 25, 26]. These included rotation to mimic variations in the orientation of lesions in clinical images, random perspective to simulate different angles and perspectives from which images can be taken during examinations, and Gaussian blur to represent common issues of slight blur due to patient movement or focus inaccuracies in dermoscopic devices. These augmentation strategies not only help prevent overfitting, but also enhance the model’s ability to generalise to new, unseen images by simulating a variety of imaging conditions in real-world scenarios.
3.2 Clinically-inspired XAI: BCC diagnosis with additional label information